### v1

@@ 3 Tensor completion via monotone inclusion
In this work, we utilize the MIP framework to address the tensor completion problem with multi-priors, constraining Eq. Eq. 7 to the tensor space $\mathbb{V}$:
(9) $\text{find }\mathcal{X}\in\mathbb{V}\text{ such that }0\in\operatorname{A}\mathcal{X}+\operatorname{B}\mathcal{X}+\operatorname{C}\mathcal{X}.$
@@
Within this formulation, any convex low-rank prior can be seamlessly incorporated by setting the corresponding operator as its subdifferential, according to Lemma 2.12. To enhance completion performance from partial observations, we extend the state-of-the-art tensor correlated total variation (t-CTV) prior to a more general and flexible form, termed generalized t-CTV (GTCTV), as detailed in Section 3.1. To further preserve fine local details, we integrate deep pseudo-contractive (DPC) denoisers. The subsequent sections introduce the formulations of GTCTV and DPC, followed by the proposed monotone-inclusion-based tensor completion model and its corresponding algorithm.
3.1 Generalized tensor correlated total variation
@@
Definition 3.1 (tensor $f$-penalty).
@@
$\|\mathcal{A}\|_{f,\mathfrak{L}}=\frac{1}{l^{2}}\sum_{i_{3}\dots i_{N}}\left\|\mathcal{A}^{(i_{3}\dots i_{N})}_{\mathfrak{L}}\right\|_{f}=\frac{1}{l^{2}}\sum_{i_{3}\dots i_{N}}\sum_{j=1}^{r}f\left(\sigma_{j}\left(\mathcal{A}^{(i_{3}\dots i_{N})}_{\mathfrak{L}}\right)\right)$,
@@
Definition 3.2 (Generalized Tensor Correlated Total Variation (GTCTV)).
Let $\mathcal{A}\in\mathbb{V}$, $\Gamma\subseteq\{1,2,\dots,N\}$ be a predefined set of gradient directions, and let $\gamma=\sharp\{\Gamma\}$ denote its cardinality, with $\nabla_{d}$ be the linear gradient operator along the $d$-th mode defined in Table 1. Then the GTCTV of $\mathcal{A}$ with respect to a regularizer $f$ and transform $\mathfrak{L}$ is
(10) $\|\mathcal{A}\|_{\operatorname{GTCTV}}:=\frac{1}{\gamma}\sum_{d\in\Gamma}\left\|\nabla_{d}\mathcal{A}\right\|_{f,\mathfrak{L}}.$
We remark that when $f(x)=|x|$, the proposed GTCTV prior reduces to the original t-CTV. To facilitate its incorporation into the MIP formulation in Eq. Eq. 9, we further provide a rigorous analysis of the weak convexity of GTCTV for a general $\mu$-weakly convex function $f$, as established in Lemma 3.3.
Lemma 3.3 (Proof in subsection SM1.2 of the supplement).
@@
3.2 Deep pseudo-contractive denoisers
@@
To address this, Wei et al. [50] introduced DPC denoisers, employing a loss function to enforce approximate pseudo-contractive properties. This approach, based on less stringent assumptions, improves PnP-based image restoration quality without imposing architectural restrictions on the network. The weaker property in Eq. Eq. 3 imply fewer constraint on the operator. As noted in [50], when the denoiser $\operatorname{D}_{\sigma}$ satisfies weaker assumption, it empirically exhibits improved denoising performance, enhancing the efficacy of PnP iterative frameworks for image restoration.
@@
$\|\operatorname{D}_{\sigma}\left(\mathcal{X}\right)-\operatorname{D}_{\sigma}\left(\mathcal{Y}\right)\|_{F}^{2}=\sum_{i_{4},\dots,i_{N}}\left\|\operatorname{D}_{\sigma}\left(\mathcal{X}^{(:,i_{4}\dots i_{N})}\right)-\operatorname{D}_{\sigma}\left(\mathcal{Y}^{(:,i_{4}\dots i_{N})}\right)\right\|_{F}^{2}$
@@
Thus, $\operatorname{D}_{\sigma}$ is $k$-SPC on $\mathbb{V}$ by Definition 2.8. The link between pseudo-contractivity and $\beta$-cocoercivity is formalized in Lemma 3.4. This result is crucial for formulating our tensor completion model within the MIP framework in Eq. Eq. 7.
Lemma 3.4 (Proof in subsection SM1.3 of the supplement).
@@
$\langle\alpha\operatorname{A}x-\alpha\operatorname{A}y,x-y\rangle=\alpha\langle\operatorname{A}x-\operatorname{A}y,x-y\rangle\geq\alpha\beta_{0}\|\operatorname{A}x-\operatorname{A}y\|^{2}=\frac{\beta_{0}}{\alpha}\|\alpha\operatorname{A}x-\alpha\operatorname{A}y\|^{2}.$
@@
3.3 Tensor completion via MIP
Drawing upon the aforementioned priors, we address tensor completion within the MIP framework of Eq. Eq. 9 by setting
(11) $\operatorname{A}=\partial\delta_{\mathcal{Y},\Omega},\,\operatorname{B}=\partial\left(\left\|\cdot\right\|_{\text{GTCTV}}+2\mu\left\|\cdot\right\|_{F}^{2}\right),\,\operatorname{C}=\alpha\left(\operatorname{Id}-\operatorname{D}_{\sigma}\right),$
where $\mathcal{Y}\in\mathbb{V}$ is the observed tensor with index set $\Omega$, and $\delta_{\mathcal{Y},\Omega}$ is the indicator function enforcing consistency:
Here, $\mu$ is the weak convexity parameter of the base function $f$, and the GTCTV prior is $4\mu$-weakly convex, as shown in Lemma 3.3. Additionally, $\alpha>0$ is a weighting factor, and $\operatorname{D}_{\sigma}$ is a deep $k$-SPC Gaussian denoiser with denoising strength $\sigma$. This yields the following concrete tensor completion model:
(12) $\text{find }\mathcal{X}\in\mathbb{V}\text{, s.t. }0\in\partial\delta_{\mathcal{Y},\Omega}\left(\mathcal{X}\right)+\partial\left(\left\|\mathcal{X}\right\|_{\text{GTCTV}}+2\mu\left\|\mathcal{X}\right\|_{F}^{2}\right)+\alpha\left(\operatorname{Id}-\operatorname{D}_{\sigma}\right)\left(\mathcal{X}\right).$
@@ 3.3 Tensor completion via MIP
Moreover, assuming $R_{\text{global}}\left(\mathcal{X}\right)=\left\|\mathcal{X}\right\|_{\text{GTCTV}}+2\mu\left\|\mathcal{X}\right\|_{F}^{2}$ and $R_{\text{local}}$ is an implicit function whose gradient is $\operatorname{Id}-\operatorname{D}_{\sigma}$, the MIP in Eq. Eq. 12 can be interpreted as the first-order optimality condition of the classical regularized optimization for tensor completion:
(13) $\min_{\mathcal{X}\in\mathbb{V}}\delta_{\mathcal{Y},\Omega}\left(\mathcal{X}\right)+R_{\text{global}}\left(\mathcal{X}\right)+R_{\text{local}}\left(\mathcal{X}\right).$
However, to the best of our knowledge, designing iterative schemes that solve Eq. Eq. 13 with guaranteed global convergence remains a significant challenge, highlighting the robustness and appeal of the monotone inclusion framework. Moreover, Wei et al. [51] show that a well-defined $R_{\text{local}}$ exists only if the deep denoiser is conservative. Imposing such a requirement introduces additional constraints on the denoiser, which may limit its effectiveness and ultimately degrade overall recovery performance.
3.4 The proposed algorithm: GTCTV-DPC
We employ the DYS method in Algorithm 1 to solve the proposed MIP Eq. Eq. 12. We begin by outlining the computation of the relevant resolvent operators with stepsize $\tau$, followed by the concrete algorithm.
3.4.1 Resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$
Since $\delta_{\mathcal{Y},\Omega}$ is convex, $\partial\delta_{\mathcal{Y},\Omega}$ is maximally monotone by Lemma 2.12. Consequently, the resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$ is the proximal operator of $\delta_{\mathcal{Y},\Omega}$:
(14) $\operatorname{J}_{\tau\partial\delta_{\mathcal{Y},\Omega}}\left(\mathcal{X}\right)=\operatorname{Prox}_{\tau\delta_{\mathcal{Y},\Omega}}\left(\mathcal{X}\right)=\mathcal{P}_{\Omega}\left(\mathcal{Y}\right)+\mathcal{P}_{\Omega^{\perp}}\left(\mathcal{X}\right),$
@@ 3.4.1 Resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$
3.4.2 Resolvent of $\tau\partial\left(\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2}\right)$
Given the convexity of $\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2}$ as proven in Lemma 3.3, the resolvent operator reduces to the proximal operator associated with $\tau\left(\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2}\right)$:
@@
$\min_{\begin{subarray}{c}\mathcal{M}\in\mathbb{V},\\ \mathcal{G}_{d},d\in\Gamma\end{subarray}}\frac{1}{\gamma}\sum_{d\in\Gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+2\mu\left\|\mathcal{M}\right\|_{F}^{2}+\frac{1}{2\tau}\left\|\mathcal{M}-\mathcal{X}\right\|_{F}^{2}\text{ s.t. }\mathcal{G}_{d}=\nabla_{d}\mathcal{M},d\in\Gamma.$
@@
(15) $\mathcal{L}\left(\mathcal{M},\left\{\mathcal{G}_{d},d\in\Gamma\right\},\left\{\mathcal{B}_{d},d\in\Gamma\right\}\right)$
$=$ $\sum_{d\in\Gamma}\left(\frac{1}{\gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+\frac{\rho_{t}}{2}\left\|\nabla_{d}\mathcal{M}-\mathcal{G}_{d}+\frac{\mathcal{B}_{d}}{\rho_{t}}\right\|_{F}^{2}\right)+2\mu\left\|\mathcal{M}\right\|_{F}^{2}+\frac{1}{2\tau}\left\|\mathcal{M}-\mathcal{X}\right\|_{F}^{2},$
@@
$\bullet$ Updating $\mathcal{M}^{t+1}$: Following [49, 48], we compute the derivative of Eq. Eq. 15 with respect to $\mathcal{M}$:
(16) $\left(\tau\rho_{t}\sum_{d\in\Gamma}\nabla_{d}^{\top}\nabla_{d}+(4\tau\mu+1)\operatorname{Id}\right)\mathcal{M}=\tau\sum_{d\in\Gamma}\nabla_{d}^{\top}(\rho_{t}\mathcal{G}_{d}-\mathcal{B}_{d})+\mathcal{X}.$
Applying multi-dimensional FFT to solve Eq. Eq. 16 yields the optimal solution for $\mathcal{M}^{t+1}$:
(17) $\mathcal{M}^{t+1}=\operatorname{F}^{-1}\left(\frac{\operatorname{F}\left(\mathcal{X}\right)+\tau\sum_{d\in\Gamma}\operatorname{F}\left(\mathcal{D}_{d}\right)^{*}\odot\operatorname{F}\left(\rho_{t}\mathcal{G}_{d}^{t}-\mathcal{B}_{d}^{t}\right)}{(4\tau\mu+1)\mathbf{1}+\tau\rho_{t}\sum_{d\in\Gamma}\operatorname{F}\left(\mathcal{D}_{d}\right)^{*}\odot\operatorname{F}\left(\mathcal{D}_{d}\right)}\right),$
@@
$\bullet$ Updating $\mathcal{G}_{d}^{t+1}$: For each $d\in\Gamma$, isolating the terms involving $\mathcal{G}_{d}$ in Eq. Eq. 15 results in the subproblem:
(18) $\mathcal{G}_{d}^{t+1}=\arg\min\limits_{\mathcal{G}_{d}\in\mathbb{V}}\frac{1}{\gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+\frac{\rho_{t}}{2}\left\|\mathcal{G}_{d}-\left(\nabla_{d}\mathcal{M}^{t+1}+\frac{\mathcal{B}_{d}^{t}}{\rho_{t}}\right)\right\|_{F}^{2}.$
To solve the subproblem in Eq. Eq. 18, we provide the proximal solution for the tensor $f$-penalty in Lemma 3.6.
Lemma 3.6 (Proof in subsection SM1.4 of the supplement).
Given $\|\cdot\|_{f,\mathfrak{L}}$ as defined in Definition 3.1, and a tensor $\mathcal{T}\in\mathbb{V}$ with t-SVD $\mathcal{T}=\mathcal{U}*_{\mathfrak{L}}\mathcal{S}*_{\mathfrak{L}}\mathcal{V}^{*}$, the solution to the proximal problem
@@
Applying Lemma 3.6 to Eq. Eq. 18, we obtain:
(19) $\mathcal{G}_{d}^{t+1}=\operatorname{t-SVF}_{\frac{1}{\gamma\rho_{t}}f}\left(\nabla_{d}\mathcal{M}^{t+1}+\frac{\mathcal{B}_{d}^{t}}{\rho_{t}}\right).$
Figure 2: Iterative flowchart of GTCTV-DPC, illustrated using a color video example. In the flowchart, the symbol $\bigoplus$ represents the operation in Eq. Eq. 14, and $\operatorname{F}^{-1}$ denotes the inverse multi-dimensional FFT in Eq. Eq. 17.
3.4.3 GTCTV-DPC

### v2

@@ III Method
In this work, we utilize the MIP framework to address the tensor completion problem with multi-priors, constraining Eq. (6) to the tensor space $\mathbb{V}$:
$\text{find }\mathcal{X}\in\mathbb{V}\text{ such that }0\in\operatorname{A}\mathcal{X}+\operatorname{B}\mathcal{X}+\operatorname{C}\mathcal{X}.$ (7)
According to Lemma 2, the operators $\operatorname{A}$ and $\operatorname{B}$ can be defined as the subdifferentials of convex functions on $\mathbb{V}$. Typically, $\operatorname{A}$ corresponds to the subdifferential of a convex data fidelity term. In the tensor completion setting, we define $\operatorname{A}=\partial\delta_{\mathcal{Y},\Omega}$, where $\mathcal{Y}\in\mathbb{V}$ denotes the observed tensor with index set $\Omega$, and $\delta_{\mathcal{Y},\Omega}$ is the indicator function enforcing data consistency:
@@
Within the formulation of Eq. (7), any convex low-rank prior can be seamlessly integrated by assigning its subdifferential to $\operatorname{B}$. To further improve completion quality under highly undersampled observations, we extend the state-of-the-art tensor correlated total variation (t-CTV) prior to a more general and flexible form, termed the generalized t-CTV (GTCTV), and set $\operatorname{B}$ accordingly in our model. To better preserve fine local details, we additionally incorporate deep pseudo-contractive (DPC) denoisers as $\operatorname{C}$. The following sections present the formulations of GTCTV and DPC, followed by the proposed monotone-inclusion-based tensor completion model and its corresponding algorithm, along with a rigorous global convergence analysis.
III-A Generalized tensor correlated total variation
@@
Definition 8 (tensor $f$-penalty).
@@
$\|\mathcal{A}\|_{f,\mathfrak{L}}$ $=\frac{1}{l^{2}}\sum_{i_{3}\dots i_{N}}\left\|\mathcal{A}^{(i_{3}\dots i_{N})}_{\mathfrak{L}}\right\|_{f}$
$=\frac{1}{l^{2}}\sum_{i_{3}\dots i_{N}}\sum_{j=1}^{r}f\left(\sigma_{j}\left(\mathcal{A}^{(i_{3}\dots i_{N})}_{\mathfrak{L}}\right)\right)$,
@@
Definition 9 (Generalized Tensor Correlated Total Variation (GTCTV)).
Let $\mathcal{A}\in\mathbb{V}$, $\Gamma\subseteq\{1,2,\dots,N\}$ be a predefined set of gradient directions, and let $\gamma=\sharp\{\Gamma\}$ denote its cardinality, with $\nabla_{d}$ be the linear gradient operator along the $d$-th mode defined in Table I. Then the GTCTV of $\mathcal{A}$ with respect to a regularizer $f$ and transform $\mathfrak{L}$ is
$\|\mathcal{A}\|_{\operatorname{GTCTV}}:=\frac{1}{\gamma}\sum_{d\in\Gamma}\left\|\nabla_{d}\mathcal{A}\right\|_{f,\mathfrak{L}}.$ (8)
We remark that when $f(x)=|x|$, the proposed GTCTV prior reduces to the original t-CTV. To facilitate its incorporation into the MIP formulation in Eq. (7), we further provide a rigorous analysis of the weak convexity of GTCTV for a general $\mu$-weakly convex function $f$, as established in Lemma 3.
Lemma 3 (Proof in section S.I.B of the supplement).
@@
III-B Deep pseudo-contractive denoisers
@@
To address this, Wei et al. [9] introduced DPC denoisers, employing a loss function to enforce approximate pseudo-contractive properties. This approach, based on less stringent assumptions, improves PnP-based image restoration quality without imposing architectural restrictions on the network. The weaker property in Eq. (3) imply fewer constraint on the operator. As noted in [9], when the denoiser $\operatorname{D}_{\sigma}$ satisfies weaker assumption, it empirically exhibits improved denoising performance, enhancing the efficacy of PnP iterative frameworks for image restoration.
@@
$\|\operatorname{D}_{\sigma}\left(\mathcal{X}\right)-\operatorname{D}_{\sigma}\left(\mathcal{Y}\right)\|_{F}^{2}$
$=$ $\sum_{i_{4},\dots,i_{N}}\left\|\operatorname{D}_{\sigma}\left(\mathcal{X}^{(:,i_{4}\dots i_{N})}\right)-\operatorname{D}_{\sigma}\left(\mathcal{Y}^{(:,i_{4}\dots i_{N})}\right)\right\|_{F}^{2}$
@@
Thus, $\operatorname{D}_{\sigma}$ is $k$-SPC on $\mathbb{V}$ by Definition 6. The link between pseudo-contractivity and $\beta$-cocoercivity is formalized in Lemma 4. This result is crucial for formulating our tensor completion model within the MIP framework in Eq. (6).
Lemma 4 (Proof in section S.I.C of the supplement).
@@
$\langle\alpha\operatorname{A}x-\alpha\operatorname{A}y,x-y\rangle=\alpha\langle\operatorname{A}x-\operatorname{A}y,x-y\rangle$
$\geq$ $\alpha\beta_{0}\|\operatorname{A}x-\operatorname{A}y\|^{2}=\frac{\beta_{0}}{\alpha}\|\alpha\operatorname{A}x-\alpha\operatorname{A}y\|^{2}.$
@@
III-C Tensor completion via MIP
Drawing upon the aforementioned priors, we address tensor completion within the MIP framework of Eq. (7) by setting
$\operatorname{A}=\partial\delta_{\mathcal{Y},\Omega},\,\operatorname{B}=\partial\left(\left\|\cdot\right\|_{\text{GTCTV}}+2\mu\left\|\cdot\right\|_{F}^{2}\right),\,\operatorname{C}=\alpha\left(\operatorname{Id}-\operatorname{D}_{\sigma}\right).$ (9)
Here, $\mu$ is the weak convexity parameter of the base function $f$, and the GTCTV prior is $4\mu$-weakly convex, as shown in Lemma 3. Additionally, $\alpha>0$ is a weighting factor, and $\operatorname{D}_{\sigma}$ is a deep $k$-SPC Gaussian denoiser with denoising strength $\sigma$. This yields the following concrete tensor completion model: find $\mathcal{X}\in\mathbb{V}$, such that
$\mathcal{O}\in$ $\partial\delta_{\mathcal{Y},\Omega}\left(\mathcal{X}\right)+\partial\left(\left\|\mathcal{X}\right\|_{\text{GTCTV}}+2\mu\left\|\mathcal{X}\right\|_{F}^{2}\right)$ (10)
$+\alpha\left(\operatorname{Id}-\operatorname{D}_{\sigma}\right)\left(\mathcal{X}\right).$
@@ III-C Tensor completion via MIP
Moreover, assuming $R_{\text{global}}\left(\mathcal{X}\right)=\left\|\mathcal{X}\right\|_{\text{GTCTV}}+2\mu\left\|\mathcal{X}\right\|_{F}^{2}$ and $R_{\text{local}}$ is an implicit function whose gradient is $\operatorname{Id}-\operatorname{D}_{\sigma}$, the MIP in Eq. (10) can be interpreted as the first-order optimality condition of the classical regularized optimization for tensor completion:
$\min_{\mathcal{X}\in\mathbb{V}}\delta_{\mathcal{Y},\Omega}\left(\mathcal{X}\right)+R_{\text{global}}\left(\mathcal{X}\right)+R_{\text{local}}\left(\mathcal{X}\right).$ (11)
However, to the best of our knowledge, designing iterative schemes that solve Eq. (11) with guaranteed global convergence remains a significant challenge, highlighting the robustness and appeal of the monotone inclusion framework. Moreover, Wei et al. [45] show that a well-defined $R_{\text{local}}$ exists only if the deep denoiser is conservative. Imposing such a requirement introduces additional constraints on the denoiser, which may limit its effectiveness and ultimately degrade overall recovery performance.
III-D The proposed algorithm: GTCTV-DPC
To solve the general MIP in Eq. (6), Davis and Yin [63] proposed the well-known Davis-Yin splitting (DYS) method. They reframe the MIP as a fixed-point problem:
$\text{find }z\in V\text{, such that }z=\operatorname{T}z,$
where $\operatorname{T}$ with stepsize $\tau$ is defined as
$\operatorname{T}:=\operatorname{Id}-\operatorname{J}_{\tau\operatorname{B}}+\operatorname{J}_{\tau\operatorname{A}}\circ(2\operatorname{J}_{\tau\operatorname{B}}-\operatorname{Id}-\tau\operatorname{C}\circ\operatorname{J}_{\tau\operatorname{B}}).$ (12)
The detailed DYS algorithm, which adopts a Krasnoselskii-Mann (KM)–type iteration [64, 65], is summarized in Algorithm 1.
Algorithm 1 DYS Algorithm for Solving the MIP in Eq. (6).
Input: $z_{0}$, $\tau$, $N_{\text{max}}$, $\{\lambda_{t}\}_{t=0}^{N_{\text{max}}}$;
for $t=0:N_{\text{max}}-1$ do
  // $z_{t+1}=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}$.
  $z_{t+1}=z_{t}+\lambda_{t}\left(x_{t+1}^{\operatorname{A}}-x_{t+1}^{\operatorname{B}}\right)$.
end for
Return $x_{t+1}^{\operatorname{A}}$.
We employ the DYS method in Algorithm 1 to solve the proposed MIP in Eq. (10). We begin by outlining the computation of the relevant resolvent operators with stepsize $\tau$, followed by the concrete algorithm.
III-D1 Resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$
Since $\delta_{\mathcal{Y},\Omega}$ is convex, $\partial\delta_{\mathcal{Y},\Omega}$ is maximally monotone by Lemma 2. Consequently, the resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$ is the proximal operator of $\delta_{\mathcal{Y},\Omega}$:
$\operatorname{J}_{\tau\partial\delta_{\mathcal{Y},\Omega}}\left(\mathcal{X}\right)=\operatorname{Prox}_{\tau\delta_{\mathcal{Y},\Omega}}\left(\mathcal{X}\right)=\mathcal{P}_{\Omega}\left(\mathcal{Y}\right)+\mathcal{P}_{\Omega^{\perp}}\left(\mathcal{X}\right),$ (13)
@@ III-D1 Resolvent of $\tau\partial\delta_{\mathcal{Y},\Omega}$
III-D2 Resolvent of $\tau\partial(\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2})$
Given the convexity of $\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2}$ as proven in Lemma 3, the resolvent operator reduces to the proximal operator associated with $\tau\left(\|\cdot\|_{\text{GTCTV}}+2\mu\|\cdot\|_{F}^{2}\right)$:
@@
$\min_{\begin{subarray}{c}\mathcal{M}\in\mathbb{V},\\ \mathcal{G}_{d},d\in\Gamma\end{subarray}}$ $\frac{1}{\gamma}\sum_{d\in\Gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+2\mu\left\|\mathcal{M}\right\|_{F}^{2}+\frac{1}{2\tau}\left\|\mathcal{M}-\mathcal{X}\right\|_{F}^{2}$
s.t. $\mathcal{G}_{d}=\nabla_{d}\mathcal{M},d\in\Gamma.$
@@
$\mathcal{L}\left(\mathcal{M},\left\{\mathcal{G}_{d},d\in\Gamma\right\},\left\{\mathcal{B}_{d},d\in\Gamma\right\}\right)$ (14)
$=$ $\sum_{d\in\Gamma}\left(\frac{1}{\gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+\frac{\rho_{t}}{2}\left\|\nabla_{d}\mathcal{M}-\mathcal{G}_{d}+\frac{\mathcal{B}_{d}}{\rho_{t}}\right\|_{F}^{2}\right)$
$+2\mu\left\|\mathcal{M}\right\|_{F}^{2}+\frac{1}{2\tau}\left\|\mathcal{M}-\mathcal{X}\right\|_{F}^{2},$
@@
$\bullet$ Updating $\mathcal{M}^{t+1}$: Following [66, 25], we compute the derivative of Eq. (14) with respect to $\mathcal{M}$:
$\left(\tau\rho_{t}\sum_{d\in\Gamma}\nabla_{d}^{\top}\nabla_{d}+(4\tau\mu+1)\operatorname{Id}\right)\mathcal{M}$ (15)
$=$ $\tau\sum_{d\in\Gamma}\nabla_{d}^{\top}(\rho_{t}\mathcal{G}_{d}-\mathcal{B}_{d})+\mathcal{X}.$
Applying multi-dimensional FFT to solve Eq. (15) yields the optimal solution for $\mathcal{M}^{t+1}$:
$\mathcal{M}^{t+1}=\operatorname{F}^{-1}\left(\frac{\operatorname{F}\left(\mathcal{X}\right)+\tau\sum_{d\in\Gamma}\operatorname{F}\left(\mathcal{D}_{d}\right)^{*}\odot\operatorname{F}\left(\rho_{t}\mathcal{G}_{d}^{t}-\mathcal{B}_{d}^{t}\right)}{(4\tau\mu+1)\mathbf{1}+\tau\rho_{t}\sum_{d\in\Gamma}\operatorname{F}\left(\mathcal{D}_{d}\right)^{*}\odot\operatorname{F}\left(\mathcal{D}_{d}\right)}\right),$ (16)
@@
$\bullet$ Updating $\mathcal{G}_{d}^{t+1}$: For each $d\in\Gamma$, isolating the terms involving $\mathcal{G}_{d}$ in Eq. (14) results in the subproblem:
$\mathcal{G}_{d}^{t+1}=\arg\min\limits_{\mathcal{G}_{d}\in\mathbb{V}}\frac{1}{\gamma}\left\|\mathcal{G}_{d}\right\|_{f,\mathfrak{L}}+\frac{\rho_{t}}{2}\left\|\mathcal{G}_{d}-\left(\nabla_{d}\mathcal{M}^{t+1}+\frac{\mathcal{B}_{d}^{t}}{\rho_{t}}\right)\right\|_{F}^{2}.$ (17)
To solve the subproblem in Eq. (17), we provide the proximal solution for the tensor $f$-penalty in Lemma 6.
Lemma 6 (Proof in section S.I.D of the supplement).
Given $\|\cdot\|_{f,\mathfrak{L}}$ as defined in Definition 8, and a tensor $\mathcal{T}\in\mathbb{V}$ with t-SVD $\mathcal{T}=\mathcal{U}*_{\mathfrak{L}}\mathcal{S}*_{\mathfrak{L}}\mathcal{V}^{*}$, the solution to the proximal problem
@@
Applying Lemma 6 to Eq. (17), we obtain:
$\mathcal{G}_{d}^{t+1}=\operatorname{t-SVF}_{\frac{1}{\gamma\rho_{t}}f}\left(\nabla_{d}\mathcal{M}^{t+1}+\frac{\mathcal{B}_{d}^{t}}{\rho_{t}}\right).$ (18)
Figure 2: Iterative flowchart of GTCTV-DPC, illustrated using a color video example. In the flowchart, the symbol $\bigoplus$ represents the operation in Eq. (13), and $\operatorname{F}^{-1}$ denotes the inverse multi-dimensional FFT in Eq. (16).
III-D3 GTCTV-DPC